OpenStudy (anonymous):

x + 3y = 5 -x + 6y = 4 solve equation this is for Algebra

OpenStudy (dmndlife24):

x + 3y =5 -x + 6y = 4 Elimination tells us to eliminate the x variables... 9y = 9 y = 1 Plugging y back into the equation... x + 3(1) = 5 x + 3 = 5 x = 2 Your answer is (2, 1)

OpenStudy (anonymous):

Thank you very much..this helped so much!

OpenStudy (dmndlife24):

No problem @cordero79

OpenStudy (anonymous):

I have one more im having problems with.. sorry im very new to this x + 3/6 = x - 6/3 equation for x

OpenStudy (dmndlife24):

Are we solving for x??

OpenStudy (anonymous):

yes

OpenStudy (dmndlife24):